OpenStudy (anonymous):

multiply the following....(5√4-5√2)(5√27+5√8)

OpenStudy (anonymous):

I am so confused by this problem

OpenStudy (anonymous):

\[ -25 \left(-2+\sqrt{2}\right) \left(2 \sqrt{2}+3 \sqrt{3}\right) \]

OpenStudy (anonymous):

{"(10-5 sqrt(2)) (10 sqrt(2)+15 sqrt(3))"}

OpenStudy (anonymous):

117.517246663902 approximately.

OpenStudy (anonymous):

(a + b)(c+d). To expand that out you would have ac+ad+bc+bd. take the elements of your equation and do the same type of thing . Your first step should look like 25(sqrt(6*27)) +25 sqrt(48) -25sqrt(54) -25sqrt(16). Then you would have to simplify all of your root expressions.

OpenStudy (anonymous):

I meant to start that response with start by looking at your equation in terms of its elements like.....

OpenStudy (anonymous):

First a times the first element in the second parenthesis and then a times the second element. Then do the same thing for b.